Express each of the following as a product of prime factors only in exponential form: 108 × 192

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The factors of 108 = 2 × 2 × 3 × 3 × 3 = 2^2× 3^3
The factors of 192 = 2 × 2 × 2 × 2 × 2 × 2 × 3 = 2^6× 3
Then,
= (2^2×3^3)×(2^6×3)
= 2^6+2 x 3^3+3 ... [∵a^m × a^n = a^m +n]
= 2^8× 3^6

Thanks